OpenStudy (superhelp101):

With the restriction x ≠ 0, Which of the following is the simplified form of 6 times x to the ninth power over 4 times x to the fourth power times the fraction 12 times x squared over 3 times x to the fifth power ? 6 over x squared 6x8 6 over x to the eighth power 6x2

OpenStudy (superhelp101):

@dumbcow @ganeshie8 can you guys help me?

OpenStudy (superhelp101):

@mathmate

ganeshie8 (ganeshie8):

can you write it using numbers/symbols ?

OpenStudy (superhelp101):

kk give me a sec ;)

OpenStudy (superhelp101):

\[(6x^9)/(4x^4) * (12x^2)/(3x^5) \] with restriction of \[x \neq0\]

ganeshie8 (ganeshie8):

\[\huge \dfrac{6x^9}{4x^4} \times \dfrac{12x^2}{3x^5}\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (superhelp101):

yes, I sorry I don't know how to use latex format ;)

ganeshie8 (ganeshie8):

Multiplying fractions is easy - you just multiply numerators separately and denominators separately

ganeshie8 (ganeshie8):

\[\huge \dfrac{6x^9\times 12x^2}{4x^4\times 3x^5} \]

OpenStudy (superhelp101):

I'm not sure how to do the restriction part

ganeshie8 (ganeshie8):

forget about the restriction, for now

ganeshie8 (ganeshie8):

it is only saying that x cannot equal 0 as it would make the expression undefined

OpenStudy (superhelp101):

you get 6x^2

ganeshie8 (ganeshie8):

Yep !and the restriction carries along : \(\large x\ne 0\)

OpenStudy (superhelp101):

what does that mean? ;)

ganeshie8 (ganeshie8):

it just means, the expression is okay for all values of \(x\), except 0

OpenStudy (superhelp101):

well I get what it's saying but I'm not sure how the 6x^2 would be affected

ganeshie8 (ganeshie8):

think of a physical situtation where the given expression models something like the cost of drinking water : \(\large f(x) = 6x^2\)

ganeshie8 (ganeshie8):

where, x = number of liters of water consumed

OpenStudy (superhelp101):

oh yeah that makes sense ;) so even if x is not equal to 0. is the answer that I got even going to change even with the restriction?

ganeshie8 (ganeshie8):

we can add the restriction : \[\large f(x) = 6x^2, ~~x\ne 0\] the function gives you cost for all values of x, except 0. because there may be a fixed water charge even if u haven't used any drinking water ?

ganeshie8 (ganeshie8):

thats one of the purposes of restrictions - they tell you for what values the function is NOT defined.

OpenStudy (superhelp101):

I see..

ganeshie8 (ganeshie8):

the cost function will not give you the cost for 0 liters of water usage, if you plugin x = 0, what would u get ?

OpenStudy (superhelp101):

0

ganeshie8 (ganeshie8):

you get \(\large f(0) = 6(0)^2 = 0\) however we need to pay a fixed charge even if we haven't used any water, right ? so we put the restriction on the function and say, the function is defined for all x except 0 : \(x \ne 0\)

ganeshie8 (ganeshie8):

this becomes more clear when u do piece wise functions :)

OpenStudy (superhelp101):

so the answer wouldn't change? it would still be 6x^2

ganeshie8 (ganeshie8):

answer is \(\large 6x^2 , ~~~x\ne 0\)

OpenStudy (superhelp101):

so the last option? ;)))

OpenStudy (mathmate):

Yes, because the restriction is part of the question and applies to all answers.